OpenStudy (anonymous):

How do you factor sin^3x + cos^3x?

myininaya (myininaya):

Do you know how to factor a sum of cubes?

myininaya (myininaya):

There is a cute formula for it.

myininaya (myininaya):

\[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

Looks cute lol thank you

myininaya (myininaya):

lol. Do you know how to use it with the trig functions?

myininaya (myininaya):

\[(\sin(x))^3+(\cos(x))^3=a^3+b^3 \]

myininaya (myininaya):

should be easy for you to determine what is a and what is b.

myininaya (myininaya):

if it isn't then let me know so i can explain further

OpenStudy (anonymous):

I've got the equation down but what to do first?

OpenStudy (anonymous):

Do I multiply what's in the parenthesis?

myininaya (myininaya):

no

myininaya (myininaya):

tell me what a is and what b is based on what I wrote above

OpenStudy (anonymous):

A is sin and b is cos

myininaya (myininaya):

\[(\sin(x))^3+(\cos(x))^3=a^3+b^3 \]

myininaya (myininaya):

ok so just go to the other side of the formula and replace a with sin(x) and replace b with cos(x) and then you are done don't do nothing more

OpenStudy (anonymous):

I did :)

myininaya (myininaya):

ok then you are done sir

myininaya (myininaya):

\[a^3+b^3=(a+b)(a^2-ab+b^2) \] everywhere we seen an a we put sin(x) everywhere we seen a b we put cos(x)

OpenStudy (anonymous):

Thanks!

myininaya (myininaya):

np :)

OpenStudy (anonymous):

I will have another question in a moment

myininaya (myininaya):

k